What mass should be hung on an elastic material of stiffness 100N/m to produce an extension of 20cm.
1
Expert's answer
2020-06-11T10:33:17-0400

By the definition of the Hooke's Law, we get:


"F = kx,"

here, "F" is the force acting on the elastic material, "k" is the stiffness of the material, "x" is the extension of the material.

From the other hand, "F = mg" (when the mass is attached), and substituting it into the first equation we get :


"mg = kx,""m = \dfrac{kx}{g},""m = \dfrac{100 \ \dfrac{N}{m} \cdot 0.2 \ m}{9.8 \ \dfrac{m}{s^2}} = 2.04 kg."

Answer:

"m = 2.04 kg."